# Illuminating Scalar Dark Matter Co-Scattering in EFT with Monophoton Signatures

### v1

@@ 2 The Model
In the above $\tilde{\Phi}=i\sigma_{2}\Phi^{\star}$ and $N^{c}=(i\gamma^{2}\gamma^{0})\overline{N}^{T}$, where $N=(N_{1},N_{2})^{T}$. Note that in the above Lagrangian flavor indices are implicit. The couplings $c_{5}$ is a $3\times 3$ matrix, $Y$ is a $3\times 2$ matrix and $c_{1,2,3}$ are $2\times 2$ matrices in flavor space. The dipole operator, $\frac{c_{3}}{\Lambda}N^{T}C^{-1}\sigma_{\mu\nu}NB^{\mu\nu}$, plays an important role in the phenomenology, in particular it leads to the decay of $N_{2}\to N_{1}\gamma$ which will give the dominant signature of the model at the LHC. This operator is relevant only when considering at least two singlet fermions, which is why we introduced two singlet fermions. Note that the renormalizable Yukawa interaction $\overline{L}\Phi N$ is forbidden due to the $Z_{2}$ symmetry. Therefore, the field $N_{i}$ can not participate in the SM neutrino mass generation at tree level via the d=5 Weinberg operator and there is no mixing between SM neutrino and $N_{i}$. After EW symmetry breaking, $c_{2}$ contribute to the mass of $N$ as $M_{N}=M_{B}+c_{2}v^{2}/\Lambda$. We redefine the couplings of the d=5 operators as, $c_{5}^{\prime}=\frac{c_{5}}{\Lambda},Y^{\prime}=\frac{Y}{\Lambda},c_{1}^{\prime}=\frac{c_{1}}{\Lambda},c_{2}^{\prime}=\frac{c_{2}}{\Lambda},c_{3}^{\prime}=\frac{c_{3}}{\Lambda}$. For simplicity, we consider all the Wilson co-efficient $c_{i}^{\prime}$ and $Y^{\prime}$ to be diagonal except for $c_{3}^{\prime}$ to be off-diagonal.
@@ 3 Dark Matter Freeze-Out
  • $\chi$ pair annihilation: $\chi\chi\to\text{SM}\leavevmode\nobreak\ \text{SM}$.
  • $\chi$ and $N_{1,2}$ co-annihilation: $\chi\leavevmode\nobreak\ N_{1,2}\to\text{SM}\leavevmode\nobreak\ \text{SM}$.
  • $N_{1,2}/\chi$ interchange: $N_{1,2}\leavevmode\nobreak\ \text{SM}\to\chi\leavevmode\nobreak\ \text{SM}$, $\chi\leavevmode\nobreak\ \nu_{e,\mu}\to N_{1,2}$ .
@@
  • $N_{1,2}$ pair annihilation: $N_{1,2}\leavevmode\nobreak\ N_{1,2}\to\gamma\gamma,\gamma Z,ZZ$.
  • $N_{1}$ and $N_{2}$ co-annihilation: $N_{1}\leavevmode\nobreak\ N_{2}\to f\bar{f},W^{+}W^{-},ZH$.
@@
For convenience and to keep with the formalism used in micrOMEGAs Alguero:2022inz we separate all particles into three sectors, namely, sector 0 corresponding to SM particles, sector $1$ comprising $N_{1}$ and $N_{2}$, and sector $2$ containing $\chi$. The particles within each sector are assumed to be in thermal equilibrium. The evolution of the dark sector particles in the early Universe is governed by the Boltzmann equations Alguero:2022inz ,
@@
$\chi$ $h$ $N_{1,2}$ $\nu_{e,\mu}$ ${Y^{{}^{\prime}}}_{11(22)}^{2}$
$\chi$ $W^{\pm}$ $N_{1,2}$ $e^{\pm}\mu^{\pm}$ ${Y^{{}^{\prime}}}_{11(22)}^{2}$ (t-channel process )
$\chi$ $\nu_{e,\mu}$ $N_{1,2}$ - ${Y^{{}^{\prime}}}_{11(22)}^{2}$ (Inverse Decay)
@@
More specifically, $\Omega h^{2}(\text{Single})$ refers to the value obtained using the standard darkOmega function in micrOMEGAs, which accounts for only one Boltzmann equation and thus incorporates solely co-annihilation processes. In contrast, $\Omega h^{2}(\text{Coupled})$ represents the result obtained via the darkOmegaN function, where the dark particles are divided into two sectors. This calculation involves solving the coupled system of Boltzmann equations 2 and 3, including all co-scattering and decay processes.
Figure 3: Feynman diagrams for the pair annihilation and co-annihilation process of $N_{1,2}$ that depend on the coupling $c_{3}^{\prime}$. Note that due to the choice $c^{\prime ij}_{3}=\epsilon^{ij}c^{\prime ij}_{3}$, we have both co-annihilation and annihilation processes between $N_{1}$ and $N_{2}$.
@@
$N_{1,2}$ $N_{1,2}$ $\gamma,Z$ $\gamma,Z$ ${c_{3}^{{}^{\prime}}}^{4}$ (t- channel process )
$N_{1}$ $N_{2}$ $f$ $\bar{f}$ ${c_{3}^{{}^{\prime}}}^{2}$ (s-channel process )
$N_{1}$ $N_{2}$ $^{+} WW^{-}$ ${c_{3}^{{}^{\prime}}}^{2}$ (s-channel process )
$N_{1}$ $N_{2}$ $Z$ $H$ ${c_{3}^{{}^{\prime}}}^{2}$ (s-channel process )
Table 2: List of all relevant annihilation and co-annihilation processes as in Fig. 3, with the dependency of their cross sections on the $c_{3}^{\prime}=c_{3}^{{}^{\prime}12}$.
Note that both sectors remain in thermal equilibrium when the Wilson coefficient ($c_{i}^{{}^{\prime}},\,Y^{{}^{\prime}}$) and the Higgs portal coupling $\lambda$ are sufficiently large. If this is the case $Y_{1}/Y_{2}=Y_{1}^{eq}/Y_{2}^{eq}$ and Eq. 23 simplify considerably and solving the coupled Boltzmann equations yields the same result as solving a single abundance equation where $Y_{1}=0$ and $Y_{2}=Y$. Hence, the usual freeze-out result is recovered and $\Delta^{1}_{\chi}=0$. Conversely, in scenarios where $Y^{{}^{\prime}},\lambda\sim\mathcal{O}(10^{-6}-10^{-10})$, $\chi$’s pair annihilation, co-annihilation, and exchange process involving $\chi$ and constituents of sector 2 become negligible, thereby simplifying the Boltzmann equations [^1] For the numerical analysis, we solve the general Boltzmann equations governing the evolution of $\chi$ and the constituents of sector 1, specifically Eqs. 2 and 3.,
@@
The dominant process in the above equation contributing to $\Gamma_{2\to 1}$ arises from $N_{1,2}$ decay while $2\to 2$ processes are subdominant. The corresponding Feynman diagrams are presented in Fig. 2. All cross-sections depend on ${Y^{{}^{\prime}}}^{2}$ as shown in Table 1. It is important to emphasize that we need at least $N_{1}$ to be in chemical equilibrium with the thermal bath through (co-)annihilation process, i.e. 1100, which depends on $c_{3}^{\prime}$ solely, to facilitate $\chi$ dilution either through co-annihilation or co-scattering. Therefore, it becomes necessary to assume $c_{3}^{\prime}>10^{-5}$ to ensure that the constituents of sector 1 remain in chemical equilibrium at $T\geq M_{N_{1,2}}$. On the contrary, due to the small coupling of $\chi$ with $N_{1,2}$ in this scenario, $\chi$ may undergo chemical decoupling at much larger $T$. The relic density of $\chi$ is set when the conversion or co-scattering processes shut off and it freezes out. Furthermore, as we focus on the conversion/co-annihilation processes, it is convenient to define the following variables, which dictate the relative mass splitting between the dark sector particles,
@@
In this work, we focus primarily on the thermal production of $\chi$ in a scenario where $c_{3}^{\prime}$ and $Y^{\prime}$ are non-zero, while fixing $\lambda=10^{-6}$ and setting all other Wilson coefficients to zero [^2] Assuming $c_{1,2}\neq 0$, $Y^{\prime}\neq 0$, and $c_{3}^{\prime}=0$, similar DM phenomenology may arise under the current assumptions of this work. A thorough exploration of additional scenarios is deferred to future work.. In this case, the value of $\delta_{2}$ determines the dominant process contributing to $N_{1}$ annihilation. When $\delta_{2}<0.5$, $N_{1}$ is primarily diluted through co-annihilation while pair annihilation is subdominant; otherwise, pair annihilation of $N_{1}$ is the dominant process Masso:2009mu ; Weiner:2012cb ; Tulin:2012uq ; Cline:2012bz . Note that the dominant co-annihilation process is $N_{1}N_{2}\to f\bar{f}$. In comparison, the process $N_{1}N_{2}\to W^{+}W^{-}$ contributes less than 1% to the total co-annihilation rate due to a cancellation between the virtual $Z$ and $\gamma$ exchange diagrams. The relevant Feynman diagrams for $N_{1}$ dilution in the early Universe are shown in Fig. 3, and the interaction rates dependence on $c_{3}^{{}^{\prime}}$ are given in Table 2.
We begin by examining two benchmark points to illustrate how the evolution of DM $\chi$ depends on the key parameters, i.e. $c_{3}^{\prime},Y^{\prime},m_{\chi}\,\text{and}\,m_{N_{1,2}}$. A detailed analysis of the allowed parameter space will be presented in Section 5. Based on the primary production mechanism of $\chi$, either through co-annihilation or co-scattering, we have considered two benchmark points. We assume $c_{3}^{\prime}=7\times 10^{-4}\leavevmode\nobreak\ \text{GeV}^{-1}$ and the other parameters are set to,
  • BP1: $M_{\chi}=100\,\text{GeV},\,M_{N_{1}}=105\,\text{GeV},\,M_{N_{2}}=115\,\text{GeV},\,{Y^{\prime}}_{11}=Y^{\prime}_{22}=1.1\times 10^{-9}\leavevmode\nobreak\ \text{GeV}^{-1}$
  • BP2: $M_{\chi}=300\,\text{GeV},\,M_{N_{1}}=315\,\text{GeV},\,M_{N_{2}}=347\,\text{GeV},\,{Y^{\prime}}_{11}=Y^{\prime}_{22}=1.2\times 10^{-9}\leavevmode\nobreak\ \text{GeV}^{-1}$
@@
It is also essential to bring to the attention of the readers that we have streamlined the calculation by assuming that $\chi$ remains in kinetic equilibrium, thus we solved the fully integrated Boltzmann equations. However, it was argued by the authors in Ref. Garny:2017rxs that dark matter may fail to maintain kinetic equilibrium along the process of conversion-driven freeze-out. A more accurate estimate of the relic abundance of the DM in the co-scattering mechanism is obtained through solving the coupled un-integrated Boltzmann equation. Nevertheless, since for our analysis the inverse decay process dominates over the $2\leftrightarrow 2$ processes, the result obtained using fully integrated Boltzmann equations are expected to differ mildly($\sim 10\%$) from those of the coupled un-integrated Boltzmann equation Brummer:2019inq .
@@ 4 Collider Constraints
Fig. 6 shows the branching ratio of $N_{2}$ as function of its mass $M_{N_{2}}$ for $M_{N_{1}},M_{\chi}=[105,100]\leavevmode\nobreak\ \rm{GeV}$, $c_{3}^{\prime}=Y_{22}^{\prime}=10^{-3}\leavevmode\nobreak\ \rm{GeV}^{-1}$. Note that there is no role of the coupling $Y_{11}^{\prime}$ in the decay of $N_{2}$. The decay $N_{2}\to N_{1}\gamma$ is the dominant channel as soon as it is kinematically accessible while $\text{Br}(N_{2}\to\chi\nu)\propto 1/{M_{N_{2}}^{2}}$ is strongly suppressed as $M_{N_{2}}$ increases. Moreover $Br(N_{2}\to N_{1}Z$) is subdominant and quickly saturates to $20\%$. For the choice of the $c_{3}^{\prime},Y_{22}^{\prime}$ to be considered (in in Table. 4)for the DM discussion, $\text{Br}(N_{2}\to\chi\nu)$ is extremely suppressed. Therefore, primary decay modes of $N_{2}$ will be $N_{2}\to N_{1}\gamma/N_{1}Z$. In Fig. 6, we show the production rate of $N_{1},N_{2}$ at the LHC. Feynman diagrams for the production processes are shown in Fig. 7. The dipole operator contributes to $pp\to N_{1}N_{2}$ and for this dominant production process the cross section scales as $(c_{3}^{\prime})^{2}$. For $N_{1}$, the only decay mode is $\chi\nu$. Hence, the dominant signature of $N_{2}$ will be mono-$\gamma$ associated with large missing momentum while mono-$Z$ is sub-dominant. The energy of $\gamma/Z$ depend directly on the mass splitting between $N_{2}$ and $N_{1}$. When $M_{N_{2}}-M_{N_{1}}$ is small, so that the photon is very soft, the mono-jet signal from $N_{1}N_{2}$ associated production with a hard-jet can become important. However, the production cross section for this process drops sharply as we demand high $p_{T}(j)$. In Fig. 6 we compare the production rate of $pp\to N_{1}N_{2}j$ (green line) for $p_{T}(j)\geq 200$ GeV with the one for $pp\to N_{1}N_{2}$ assuming a mass splitting of 5 GeV between $N_{2}$ and $N_{1}$. Note that the mono-jet signal can arise from the process $pp\to N_{i}N_{i}j$ via the coupling $c_{2}^{\prime}$, here it is irrelevant since we set $c_{2}^{\prime}=0$. The production of a DM pair via the coupling with Higgs, $pp\to h\to\chi\chi j$ can also lead to a mono-jet signal but the cross section is suppressed by the coupling $\lambda$ considered, $\lambda<10^{-6}$. We obtain that the mono-jet cross section in our model (for the parameter space of interest) is one order of magnitude smaller than the observed cross section from ATLAS search ATLAS:2021kxv . Therefore mono-$\gamma$ gives the leading signal. We consider only this process in the following.
@@
There exist an ATLAS search for events with an energetic photon and large missing transverse momentum at the 13 TeV LHC with integrated luminosity of $139\leavevmode\nobreak\ \text{fb}^{-1}$ ATLAS:2020uiq . As mentioned above, the production of $N_{1}N_{2}$ via the channel $pp\to N_{1}N_{2}$ mediated by the dipole operator ${c_{3}^{\prime}}\overline{N^{c}}\sigma_{\mu\nu}NB^{\mu\nu}$ can lead to such signal for the mass hierarchy $M_{N_{2}}>M_{N_{1}}>M_{\chi}$ such that $N_{2}\to N_{1}\gamma$ and $N_{1}$ decay invisibly, $N_{1}\to\chi\nu$ leading to $pp\to N_{1}N_{2}\to\gamma+E_{T}^{\rm miss}$. The mass splitting $M_{N_{2}}-M_{N_{1}}$ is crucial for the sensitivity of the signal as it governs the energy of the photon. Figure 8 shows the distribution of $p_{T}^{\gamma}$ for $M_{N_{2}}=[120,150,200]$ GeV and $M_{N_{1}},M_{\chi}=[105,100]$ GeV. As expected, the photon is harder for larger $M_{N_{2}}-M_{N_{1}}$. The signal regions in the ATLAS search ATLAS:2020uiq demands $E_{T}^{\gamma}>150$ GeV and upper limits at $95\%$ CL on the visible cross section are set for 7 signal regions (SRs): 4 inclusive (SRI1–SRI4) and 3 exclusive (SRE1-SRE3). The SRs are defined corresponding to different $E_{T}^{\rm miss}$ ranges, as listed in Table 3.
@@
Table 3: Upper limits at $95\%$ CL on the visible cross section $(\sigma\times A\times\epsilon)_{95}^{\rm obs}$ of the signal $pp\to\gamma+E_{T}^{\rm miss}$ for various signal regions (SRs) ATLAS:2020uiq .
Figure 9: Limit from ATLAS search ATLAS:2020uiq in the $(c_{3}^{\prime})_{UL}$ vs $M_{N_{2}}$ plane for signal region SRE1 (red) and SRI4 (black) for (a) $M_{\chi}=100$ GeV, $M_{N_{1}}=105$ GeV (b) $M_{\chi}=300$ GeV, $M_{N_{1}}=315$ GeV. In both cases $Y_{22}^{\prime}=10^{-9}\leavevmode\nobreak\ {\rm GeV}^{-1}$.
Figure 10: Upper limit on $c_{3}^{\prime}$ from the ATLAS search ATLAS:2020uiq in the $\delta_{1}-\delta_{2}$ plane for (a) $M_{\chi}=100$ GeV and (b) $M_{\chi}=300$ GeV. Here $Y_{22}^{\prime}=10^{-9}\leavevmode\nobreak\ {\rm GeV^{-1}}$.

### v2

@@ 2 The Model
In the above equation $L$, $\Phi$ and $B_{\mu\nu}$ are the SM Lepton doublet, Higgs doublet and hypercharge field strength tensor, respectively; additionally, $\tilde{\Phi}=i\sigma_{2}\Phi^{\star}$ and $N^{c}=(i\gamma^{2}\gamma^{0})\overline{N}^{T}$, where $N=(N_{1},N_{2})^{T}$. Note that in the above Lagrangian flavor indices are implicit. The couplings $c_{5}$ is a $3\times 3$ matrix, $Y$ is a $3\times 2$ matrix and $c_{1,2,3}$ are $2\times 2$ matrices in flavor space. The dipole operator, $\frac{c_{3}}{\Lambda}N^{T}C^{-1}\sigma_{\mu\nu}NB^{\mu\nu}$, plays an important role in the phenomenology, in particular it leads to the decay of $N_{2}\to N_{1}\gamma$ which will give the dominant signature of the model at the LHC. This operator is relevant only when considering at least two singlet fermions, which is why we introduced two singlet fermions. Note that the renormalizable Yukawa interaction $\overline{L}\Phi N$ is forbidden due to the $Z_{2}$ symmetry. Therefore, the field $N_{i}$ can not participate in the SM neutrino mass generation at tree level via the d=5 Weinberg operator and there is no mixing between SM neutrino and $N_{i}$. After electroweak symmetry breaking, $c_{2}$ contribute to the mass of $N$ as $M_{N}=M_{B}+c_{2}v^{2}/\Lambda$. We redefine the couplings of the d=5 operators as, $c_{5}^{\prime}=\frac{c_{5}}{\Lambda},y^{\prime}=\frac{y}{\Lambda},c_{1}^{\prime}=\frac{c_{1}}{\Lambda},c_{2}^{\prime}=\frac{c_{2}}{\Lambda},c_{3}^{\prime}=\frac{c_{3}}{\Lambda}$. For simplicity, we consider all the Wilson co-efficient $c_{i}^{\prime}$ and $y^{\prime}$ to be diagonal except for $c_{3}^{\prime}$ to be off-diagonal.
@@ 3 Dark Matter Freeze-Out
  • $\chi$ pair annihilation: $\chi\chi\to\text{SM}\penalty 10000\ \text{SM}$.
  • $\chi$ and $N_{1,2}$ co-annihilation: $\chi\penalty 10000\ N_{1,2}\to\text{SM}\penalty 10000\ \text{SM}$.
  • $N_{1,2}/\chi$ interchange: $N_{1,2}\penalty 10000\ \text{SM}\to\chi\penalty 10000\ \text{SM}$, $\chi\penalty 10000\ \nu_{e,\mu}\to N_{1,2}$ .
@@
  • $N_{1,2}$ pair annihilation: $N_{1,2}\penalty 10000\ N_{1,2}\to\gamma\gamma,\gamma Z,ZZ$.
  • $N_{1}$ and $N_{2}$ co-annihilation: $N_{1}\penalty 10000\ N_{2}\to f\bar{f},W^{+}W^{-},ZH$.
@@
For convenience and to keep with the formalism used in micrOMEGAs Alguero:2022inz we separate all particles into three sectors, namely, sector 0 corresponding to SM particles, sector $1$ comprising $N_{1}$ and $N_{2}$, and sector $2$ containing $\chi$  [^1] Here we will refer to the dark sectors either as sector 1 and sector 2 or as $N\,(=N_{1}+N_{2})$ and $\chi$, respectively. . The particles within each sector are assumed to be in thermal equilibrium. The evolution of the dark sector particles in the early Universe is governed by the Boltzmann equations Alguero:2022inz ,
@@
$\chi$ $h$ $N_{1,2}$ $\nu_{e,\mu}$ ${y^{\prime}}_{11(22)}^{2}$
$\chi$ $W^{\pm}$ $N_{1,2}$ $e^{\pm}\mu^{\pm}$ ${y^{\prime}}_{11(22)}^{2}$ (t-channel process )
$\chi$ $\nu_{e,\mu}$ $N_{1,2}$ - ${y^{\prime}}_{11(22)}^{2}$ (Inverse Decay)
@@
More specifically, $\Omega h^{2}(\text{Single})$ refers to the value obtained using the standard darkOmega function in micrOMEGAs, which accounts for only one Boltzmann equation and assumes that all particles are in thermal equilibrium in the early Universe. Thus it incorporates solely co-annihilation processes. In contrast, $\Omega h^{2}(\text{Coupled})$ represents the result obtained via the darkOmegaN function, where the dark particles are divided into two sectors that are not necessarily in thermal equilibrium with each other. This calculation involves solving the coupled system of Boltzmann equations 2 and 3, including all co-scattering and decay processes.
Figure 3: Feynman diagrams for the pair annihilation and co-annihilation process of $N_{1,2}$ that depend on the coupling $c_{3}^{\prime}$. Note that due to the choice $(c^{\prime}_{3})_{ij}=\epsilon_{ij}(c^{\prime}_{3})_{ij}$, we have both co-annihilation and annihilation processes between $N_{1}$ and $N_{2}$.
@@
$N_{1,2}$ $N_{1,2}$ $\gamma,Z$ $\gamma,Z$ ${c_{3}^{\prime}}^{4}$ (t- channel process )
$N_{1}$ $N_{2}$ $f$ $\bar{f}$ ${c_{3}^{\prime}}^{2}$ (s-channel process )
$N_{1}$ $N_{2}$ $^{+} WW^{-}$ ${c_{3}^{\prime}}^{2}$ (s-channel process )
$N_{1}$ $N_{2}$ $Z$ $H$ ${c_{3}^{\prime}}^{2}$ (s-channel process )
Table 2: List of all relevant annihilation and co-annihilation processes as in Fig. 3, with the dependency of their cross sections on the $c_{3}^{\prime}=(c_{3}^{\prime})_{12}$.
Note that both sectors remain in thermal equilibrium when the Wilson coefficient ($c_{i}^{\prime},\,y^{\prime}$) and the Higgs portal coupling $\lambda$ are sufficiently large. If this is the case $Y_{1}/Y_{2}=Y_{1}^{eq}/Y_{2}^{eq}$ and Eq. 2, 3 simplify considerably and solving the coupled Boltzmann equations yields the same result as solving a single abundance equation where $Y_{1}=0$ and $Y_{2}=Y$. Hence, the usual freeze-out result is recovered and $\Delta^{1}_{\chi}=0$. Conversely, in scenarios where $y^{\prime},\lambda\sim\mathcal{O}(10^{-6}-10^{-10})$, $\chi$’s pair annihilation, co-annihilation, and exchange process involving $\chi$ and constituents of sector 2 become negligible, thereby simplifying the Boltzmann equations [^2] For the numerical analysis, we solve the general Boltzmann equations governing the evolution of $\chi$ and the constituents of sector 1, specifically Eqs. 2 and 3.,
@@
The dominant process in the above equation contributing to $\Gamma_{2\to 1}$ arises from $N_{1,2}$ decay while $2\to 2$ processes are subdominant. The corresponding Feynman diagrams are presented in Fig. 2. All cross-sections depend on $y^{\prime 2}$ as shown in Table 1. It is important to emphasize that we need at least $N_{1}$ to be in chemical equilibrium with the thermal bath through (co-)annihilation process, i.e. 1100, which depends on $c_{3}^{\prime}$ solely, to facilitate $\chi$ dilution either through co-annihilation or co-scattering. Note that 1100 refers to processes of the type $N_{i}N_{j}\to SM,SM$, these are illustrated in Fig. 3. Therefore, it becomes necessary to assume $c_{3}^{\prime}>10^{-5}$ to ensure that the constituents of sector 1 remain in chemical equilibrium at $T\geq M_{N_{1,2}}$. On the contrary, due to the small coupling of $\chi$ with $N_{1,2}$ in this scenario, $\chi$ may undergo chemical decoupling at much larger $T$. The relic density of $\chi$ is set when the conversion or co-scattering processes shut off and it freezes out. Furthermore, as we focus on the conversion/co-annihilation processes, it is convenient to define the following variables, which dictate the relative mass splitting between the dark sector particles,
@@
In this work, we focus primarily on the thermal production of $\chi$ in a scenario where $c_{3}^{\prime}$ and $y^{\prime}$ are non-zero, while fixing $\lambda=10^{-6}$ and setting all other Wilson coefficients to zero [^3] Assuming $c_{1,2}\neq 0$, $y^{\prime}\neq 0$, and $c_{3}^{\prime}=0$, similar DM phenomenology may arise under the current assumptions of this work. A thorough exploration of additional scenarios is deferred to future work.. In this case, the value of $\delta_{2}$ determines the dominant process contributing to $N_{1}$ annihilation. When $\delta_{2}<0.5$, $N_{1}$ is primarily diluted through co-annihilation while pair annihilation is subdominant; otherwise, pair annihilation of $N_{1}$ is the dominant process Masso:2009mu ; Weiner:2012cb ; Tulin:2012uq ; Cline:2012bz . Note that the dominant co-annihilation process is $N_{1}N_{2}\to f\bar{f}$. In comparison, the process $N_{1}N_{2}\to W^{+}W^{-}$ contributes less than 1% to the total co-annihilation rate due to a cancellation between the virtual $Z$ and $\gamma$ exchange diagrams. The relevant Feynman diagrams for $N_{1}$ dilution in the early Universe are shown in Fig. 3, and the interaction rates dependence on $c_{3}^{\prime}$ are given in Table 2. It is important to bring to the attention of the readers that a wide range of DM masses is feasible in our model, spanning from sub-GeV to the TeV scale. In our work, we primarily focus on weak-scale DM mass to have a viable collider signature in the LHC, as will be discussed in the next section. In contrast, neither upcoming direct nor indirect detection experiments can provide viable signatures, since the coupling responsible for dark matter detection, $\lambda$, is set to $10^{-6}$.
We begin by examining two benchmark points to illustrate how the evolution of DM $\chi$ depends on the key parameters, i.e. $c_{3}^{\prime},\,y^{\prime},M_{\chi}\,\text{and}\,M_{N_{1,2}}$. A detailed analysis of the allowed parameter space will be presented in Section 5. Based on the primary production mechanism of $\chi$, either through co-annihilation or co-scattering, we have considered two benchmark points. We assume $c_{3}^{\prime}=7\times 10^{-4}\penalty 10000\ \text{GeV}^{-1}$ and the other parameters are set to,
  • BP1: $M_{\chi}=100\,\text{GeV},\,M_{N_{1}}=105\,\text{GeV},\,M_{N_{2}}=115\,\text{GeV},\,{y^{\prime}}_{11}=y^{\prime}_{22}=1.1\times 10^{-9}\penalty 10000\ \text{GeV}^{-1}$
  • BP2: $M_{\chi}=300\,\text{GeV},\,M_{N_{1}}=315\,\text{GeV},\,M_{N_{2}}=347\,\text{GeV},\,{y^{\prime}}_{11}=y^{\prime}_{22}=1.2\times 10^{-9}\penalty 10000\ \text{GeV}^{-1}$
@@
It is important to point out that we have streamlined the calculation by assuming that $\chi$ remains in kinetic equilibrium, thus we solved the fully integrated Boltzmann equations. In our analysis, the inverse decay process dominates over the $2\leftrightarrow 2$ interactions. Consequently, the results obtained using the fully integrated Boltzmann equations are expected to differ by less than $10\%$ from those of the coupled un-integrated Boltzmann equations Garny:2017rxs ; Alguero:2022inz . This small deviation arises because the decay and inverse decay processes remain efficient until late times, thereby re-establishing an approximate kinetic equilibrium (see Fig. 8 of Ref. Garny:2017rxs ). However, in scenarios where the inverse decay becomes subdominant, the deviation from the coupled un-integrated Boltzmann results can become significant, as the dark matter momentum distribution becomes highly distorted (i.e., deviates considerably from a thermal distribution). In such cases, the integrated Boltzmann equation may under- or overestimate the actual relic yield Brummer:2019inq , thus solving the coupled un-integrated Boltzmann equation will provide a more accurate estimate of the relic abundance of the DM. We leave this for future work.
@@ 4 Collider Constraints
Fig. 6 shows the branching ratio of $N_{2}$ as function of its mass $M_{N_{2}}$ for $M_{N_{1}},M_{\chi}=[105,100]\penalty 10000\ \rm{GeV}$, $c_{3}^{\prime}=y_{22}^{\prime}=10^{-3}\penalty 10000\ \rm{GeV}^{-1}$. Note that there is no role of the coupling $Y_{11}^{\prime}$ in the decay of $N_{2}$. The decay $N_{2}\to N_{1}\gamma$ is the dominant channel as soon as it is kinematically accessible while $\text{Br}(N_{2}\to\chi\nu)\propto 1/{M_{N_{2}}^{2}}$ is strongly suppressed as $M_{N_{2}}$ increases. Moreover $Br(N_{2}\to N_{1}Z$) is subdominant and quickly saturates to $20\%$. For the choice of the $c_{3}^{\prime},y_{22}^{\prime}$ to be considered (in Table. 4)for the DM discussion, $\text{Br}(N_{2}\to\chi\nu)$ is extremely suppressed. Therefore, primary decay modes of $N_{2}$ will be $N_{2}\to N_{1}\gamma/N_{1}Z$. In Fig. 6, we show the production rate of $N_{1},N_{2}$ at the LHC. Feynman diagrams for the production processes are shown in Fig. 7. The dipole operator contributes to $pp\to N_{1}N_{2}$ and for this dominant production process the cross section scales as $(c_{3}^{\prime})^{2}$. For $N_{1}$, the only decay mode is $\chi\nu$. Hence, the dominant signature of $N_{2}$ will be mono-$\gamma$ associated with large missing momentum while mono-$Z$ is sub-dominant. The energy of $\gamma/Z$ depend directly on the mass splitting between $N_{2}$ and $N_{1}$. When $M_{N_{2}}-M_{N_{1}}$ is small, so that the photon is very soft, the mono-jet signal from $N_{1}N_{2}$ associated production with a hard-jet can become important. However, the production cross section for this process drops sharply as we demand high $p_{T}(j)$. In Fig. 6 we compare the production rate of $pp\to N_{1}N_{2}j$ (green line) for $p_{T}(j)\geq 200$ GeV with the one for $pp\to N_{1}N_{2}$ assuming a mass splitting of 5 GeV between $N_{2}$ and $N_{1}$. Note that the mono-jet signal can arise from the process $pp\to N_{i}N_{i}j$ via the coupling $c_{2}^{\prime}$, here it is irrelevant since we set $c_{2}^{\prime}=0$. The production of a DM pair via the coupling with Higgs, $pp\to h\to\chi\chi j$ can also lead to a mono-jet signal but the cross section is suppressed by the coupling $\lambda$ considered, $\lambda<10^{-6}$. We obtain that the mono-jet cross section in our model (for the parameter space of interest) is one order of magnitude smaller than the observed cross section from ATLAS search ATLAS:2021kxv . Therefore mono-$\gamma$ gives the leading signal. We consider only this process in the following.
@@
There exist an ATLAS search for events with an energetic photon and large missing transverse momentum at the 13 TeV LHC with integrated luminosity of $139\penalty 10000\ \text{fb}^{-1}$ ATLAS:2020uiq . As mentioned above, the production of $N_{1}N_{2}$ via the channel $pp\to N_{1}N_{2}$ mediated by the dipole operator ${c_{3}^{\prime}}\overline{N^{c}}\sigma_{\mu\nu}NB^{\mu\nu}$ can lead to such signal for the mass hierarchy $M_{N_{2}}>M_{N_{1}}>M_{\chi}$ such that $N_{2}\to N_{1}\gamma$ and $N_{1}$ decay invisibly, $N_{1}\to\chi\nu$ leading to $pp\to N_{1}N_{2}\to\gamma+E_{T}^{\rm miss}$. The mass splitting $M_{N_{2}}-M_{N_{1}}$ is crucial for the sensitivity of the signal as it governs the energy of the photon. Figure 8 shows the distribution of $p_{T}^{\gamma}$ for $M_{N_{2}}=[120,150,200]$ GeV and $M_{N_{1}},M_{\chi}=[105,100]$ GeV. As expected, the photon is harder for larger $M_{N_{2}}-M_{N_{1}}$. The signal regions in the ATLAS search ATLAS:2020uiq demands $E_{T}^{\gamma}>150$ GeV and upper limits at $95\%$ CL on the visible cross section are set for 7 signal regions (SRs): 4 inclusive (SRI1–SRI4) and 3 exclusive (SRE1-SRE3). The SRs are defined corresponding to different $E_{T}^{\rm miss}$ ranges, as listed in Table 3.
@@
Table 3: Upper limits at $95\%$ CL on the visible cross section $(\sigma\times A\times\epsilon)_{95}^{\rm obs}$ of the signal $pp\to\gamma+E_{T}^{\rm miss}$ for various signal regions (SRs), where $A$ is the acceptance and $\epsilon$ is the efficiency ATLAS:2020uiq .
Figure 9: Limit from ATLAS search ATLAS:2020uiq in the $(c_{3}^{\prime})_{UL}$ vs $M_{N_{2}}$ plane for signal region SRE1 (red) and SRI4 (black) for (a) $M_{\chi}=100$ GeV, $M_{N_{1}}=105$ GeV (b) $M_{\chi}=300$ GeV, $M_{N_{1}}=315$ GeV. In both cases $y_{22}^{\prime}=10^{-9}\penalty 10000\ {\rm GeV}^{-1}$.
Figure 10: Upper limit on $c_{3}^{\prime}$ from the ATLAS search ATLAS:2020uiq in the $\delta_{1}-\delta_{2}$ plane for (a) $M_{\chi}=100$ GeV and (b) $M_{\chi}=300$ GeV. Here $y_{22}^{\prime}=10^{-9}\penalty 10000\ {\rm GeV^{-1}}$.